a line is such that its segment between the straight lines 5x-y-4=0 and 3x+4y-4=0 is bisected at the point (1,5).obtain the equation?

Grade:11

4 Answers

bill king
11 years ago

Assume the pts. where they cut as (a,b) and (c,d).

Then put them in the equation of the respective lines.

Als: a+c=2;b+d=10

Thus you will get

83x-35y+92=0

Thank you

Ayushi Agrawal
11 years ago

wats this a+c=2 and b+d=10...........

how??????????

bill king
11 years ago

Since (1,5) is the mid point of (a,b) and (c,d)

therefore (a+c)/2=1

that is why a+c=2

(b+d)/2=5

that is why b+d=10.

Thank you

harshit
10 years ago
how do we put these in the eq